### v1

@@ 4.2 Dependence on the strength of the self-interactions
Figure 4: Similar to Fig. 3, we show the position and velocity of the BH particle as a function of time. The results are shown for several interaction strengths of the DM, indicated by the opacity of the lines. All simulations are for the large box with the high resolution of $N_{\mathrm{DM}}=5\times 10^{6}$ particles. For the simulation of gaseous matter, we used the same number of SPH particles.
@@
Given that the dynamical friction force differs between the collisionless and the gaseous case as described in Section 2, we expect that self-interactions can have an impact on the strength of the dynamical friction force. To study this in detail, we run simulations with SIDM. In Fig. 4, we show the results for fSIDM with different values for the momentum transfer cross-section. We also include the collisionless and gaseous case. It is visible that the self-interactions can enhance or reduce the dynamical friction, depending on the strength of the cross-section. For the gaseous regime, we expect the dynamical friction to be enhanced above the sound speed and reduced below. Accordingly, we find a larger velocity of the BH at the end of the simulation, when the BH velocity has fallen well below the sound speed of the DM.
@@
Figure 5: We show the surface density (top row) and the velocity dispersion (bottom row) for our test simulations at a time of $t=0.49\,\mathrm{yr}$. The panels display the different types of matter we investigate, with increasing collisionality from the left to the right. The density and one dimensional velocity dispersion are indicated following a logarithmic colour scale spanning a factor of two. For all panels of the same quantity, the same colour scale is used. In addition, we indicate the position of the BH particle with a red dot. In the online ancillary files, we provide the time evolution of the surface density as a video.
@@
A detailed investigation of the dynamic friction force will follow. But before, we take a look at the DM density perturbed by the BH. For the same simulations as shown in Fig. 4, we show the surface density of the DM in the top row of Fig. 5. Here, we can see that the shape and density of the wake depend on the strength of the self-interactions. We find that for $\sigma_{\mathrm{V}}/m\in\{3\times 10^{-8},3\times 10^{-7}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$ the densities in the wake are higher and at the same time the deceleration is stronger. For these small cross-sections, the scattering leads to more low-velocity particles in the density wake compared to CDM effectively enhancing the deceleration of the BH.
@@
In contrast, for the cases of even stronger self-interactions, we find that the densities in the wake are decreasing and the deceleration of the BH is weaker. But also note that the shape of the density wake changes for an increasing self-interaction strength, impacting the dynamical friction force as well. The reduced dynamical friction for the strong cross-sections is in line with the expectation of a weaker deceleration in the gaseous case for the subsonic regime.
@@
Figure 6: We show the dynamical friction force as a function of the Mach number for our simulations. This plot is analogue to Fig. 2, which gives the analytic description of the dynamical friction force. Different cross-sections are shown, indicated by the Knudsen number. Note, $\mathrm{Kn}=\infty$ corresponds to a collisionless medium and $\mathrm{Kn}=0.0$ to a gaseous medium.
For a more detailed analysis of the dynamical friction force and its dependence on the cross-section, we compute the force and show it as a function of the velocity (Fig. 6). We note, the strongest SIDM cross-section shown is $\sigma_{\mathrm{V}}/m=3\times 10^{-6}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. Given our resolution, we are not able to faithfully simulate a cross-section as large as $\sigma_{\mathrm{V}}/m=3\times 10^{-5}\mathrm{cm}^{2}\mathrm{g}^{-1}$ because the mean free path becomes much smaller than the size of the kernel used for the self-interactions (see Appendix A and Fischer et al., 2024a).
We specify the strength of the cross-section in terms of the Knudsen number (Eq. 10). The collisionless case is indicated by $\mathrm{Kn}=\infty$ and the gaseous case by $\mathrm{Kn}=0.0$. In Fig. 6 we can see that the dynamical friction is initially increasing (starting at the highest velocity) and decreases for later times of the simulation. At the beginning of the simulation, where the BH moves with a supersonic velocity, the dynamical friction is the smallest in the collisionless case. At later stages, it depends on the cross-section whether the self-interactions enhance or reduce the deceleration of the BH. In Fig.6, we plot the dynamical friction as a function of the velocity. This shows that the decelerating force reaches its maximum roughly at the sound speed, as expected. Interestingly, the self-interacting run with an intermediate cross-section ($\mathrm{Kn}=1.35\times 10^{-1}$) shows the strongest dynamical friction, stronger than in the collisionless or gaseous regime. In principle, the figure is built analogously to Fig. 2, but a direct comparison is limited by the fact that our simulation setup may not fulfill the assumptions of the Ostriker formula for the dynamical friction in the gaseous case (see Section 2.2). This is for example illustrated by Fig. 3, where we have seen that the exact results depend on the choice of the setup, in particular on the box size.
@@ 4.3 Properties of the wake
Figure 7: The density of the wake, i.e., the mean density about the point with the highest density, is shown as a function of the velocity of the perturbing BH. The simulation results of the different cross-sections are indicated by their Knudsen number.
Figure 8: The one-dimensional velocity dispersion of the wake, i.e., the velocity dispersion in the vicinity of the point with the highest density, is shown as a function of the velocity of the perturber. We note that the velocity dispersion is computed after the bulk motion has been subtracted. The simulation results for different cross-sections are shown, indicated by the Knudsen number.
Figure 9: The mean velocity of the wake in the direction of the motion of the BH, i.e., the mean velocity in the vicinity of the point with the highest density, is shown. The results for different cross-sections are shown as a function of the velocity of the perturber. They are indicated by their corresponding Knudsen number.
Figure 10: The DM mass gravitational bound to the perturbing BH is shown as a function of the velocity of the perturber. We indicate the different cross-sections by their Knudsen number.
To provide more insight into the increased dynamical friction for mildly collisional media, we compute further quantities such as the density of the wake, its velocity dispersion, and bulk motion as well as the DM mass gravitational bound to the BH.
@@
Furthermore, we show the velocity dispersion in the density wake in Fig. 8. We use the same volume as for the density computation. The velocity dispersion is computed after subtracting the mean velocity of the particles. The velocity dispersion in the wake increases compared to the initial conditions and is larger for stronger DM self-interactions. Interestingly, the cross-section with $\mathrm{Kn}=1.35\times 10^{-1}$, which gives the strongest dynamical friction, shows a substantially lower velocity dispersion compared to the larger cross-sections.
The bulk motion of the density wake is displayed in Fig. 9. The stronger the self-interactions the more efficiently the BH drags the DM along its trajectory, except for the largest cross-section for which the bulk velocity is again smaller. This is in line with the strength of the dynamical friction force found in Fig. 6, i.e., when the DM is dragged along by the BH, it can more efficiently decelerate the BH.
Last, we compute the DM mass that is gravitationally bound to the BH. To do so, we compare the relative velocity of a DM particle and the BH to the escape velocity,
@@ 4.4 Angular dependence
The angular dependence of the self-interaction cross-section can in principle lead to qualitative differences in the evolution of astrophysical systems. This is for example well known for merging DM halos (e.g., Fischer et al., 2023; Sabarish et al., 2024) or the abundance of satellites (Fischer et al., 2022). However, for systems that are rather relaxed, the angular dependence plays a minor role only and the effect of self-interactions with different angular dependencies can be mapped onto each other using the viscosity cross-section $\sigma_{\mathrm{V}}$ (Eq. 11, see also Yang & Yu, 2022).
To test whether this mapping also works for our setup to study dynamical friction, we run simulations with a very anisotropic cross-section (using the fSIDM framework) and an isotropic cross-section. The cross-sections are chosen to have the same value in terms of $\sigma_{\mathrm{V}}/m$ or $\sigma_{\mathrm{T}}/m$ to test the matching with the momentum transfer cross-section (see Eq. 12) as well.
In Fig. 11, we compare the velocity of the BH for the different angular dependencies. We compute the ratio of the change in BH velocities for simulations with a matched isotropic cross-section to the fiducial fSIDM simulations. The two panels in Fig. 11 display the ratio of the velocities as a function of time for a weaker fSIDM cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ (top panel) and a stronger one of $\sigma_{\mathrm{V}}/m=3\times 10^{-7}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ (bottom panel). For a perfect matching, we would expect the ratio to be unity. As we can see, the viscosity cross-section (Eq. 11) provides a better match than the transfer cross-section (Eq. 12). This is the case for the weaker and stronger cross-section that we have simulated to test this.
Figure 11: We compute the ratio between the velocity change, $\Delta v=v(0)-v(t)$, due to dynamical friction of an fSIDM cross-section and the transfer (blue) / viscosity (orange) matched isotropic cross-sections. The upper panel is for an fSIDM viscosity cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ and the lower panel is for $\sigma_{\mathrm{V}}/m=3\times 10^{-7}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$.
@@ 4.5 Velocity dependence
The effective cross-section $\sigma_{\mathrm{eff}}$ (introduced by Yang & Yu, 2022) allows to map cross-sections with different velocity dependencies onto each other. It integrates the cross-section with its velocity dependence over the assumed Maxwell-Boltzmann distribution of the scattering velocities, applying a weighting factor of $v^{5}$. The effective cross-section is given by
@@
Note that this is very similar to the matching procedure proposed by Yang et al. (2023); see also the work by Outmezguine et al. (2023). It has been shown that this works fairly well for isolated halos and, in turn, velocity-dependent cross-sections can be approximately described with constant cross-sections. However, it is known that this does not work for systems that involve multiple velocity scales, such as merging galaxy clusters (Sabarish et al., 2024) or halos containing substructure (Fischer et al., 2024b).
@@
Figure 12: The ratio of the velocity change, $\Delta v=v(0)-v(t)$, due to dynamical friction acting on the BH for a velocity-dependent and a velocity-independent (constant) cross-section is shown. All cross-sections were matched using the effective cross-section (Eq. 15), i.e., all cross-sections correspond to $\sigma_{\mathrm{eff}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. The upper panel displays runs with frequent self-interactions, and the lower panel is for isotropic scattering.
@@ 5 Discussion
For binary BHs, we find that the motion of the secondary BH through the DM spike would typically happen at Mach numbers larger than unity, i.e., in the supersonic regime. Moreover, the Mach number does only very slightly depend on the radius, as the mass of the DM spike in the relevant radial range is typically much smaller than the mass of the primary BH. Also, the spike index does not play a significant role in the motion of the secondary BH being supersonic. But more importantly, a non-zero eccentricity could lead to a subsonic phase about the apocentre of the orbit of the secondary BH. This implies that the supersonic regime where the dynamical friction is enhanced by self-interactions plays an important role and would lead to a faster decay of the orbit compared to a collisionless medium with the same density.
To study the dynamical friction with self-interactions being present, simulations are required. We used an idealized setup with a periodic box hosting initially a constant density. We have to note that this setup comes with drawbacks that limit the conclusion we can draw. In particular, it is difficult to create a stable setup where the box size is small enough such that the velocity dispersion of the DM counteracts the gravitational collapse. This effectively limits the maximal box size that can be studied, which also has an impact on the dynamical friction. This is because the box size is related to the maximum impact parameter $r_{\mathrm{max}}$ (see Eq. 7). Moreover, the setup does not always fulfill the assumptions of analytic formulas discussed in section 2.2. Also, we cannot exclude that perturbations arising from the supersonic phase still affect the trajectory of the BH in the subsonic phase. Nevertheless, we were still able to make qualitative statements and gain new insights into the dynamical friction. Especially for this purpose, we tuned the setup to give a relatively larger impact on the BH velocity and make it easier to obtain qualitative differences being much larger than the numerical errors.
@@ 6 Conclusions
In this paper, we have reviewed analytic estimates for the dynamical friction force in the collisionless and gaseous regime, and discussed the two regimes in the context of binary BH mergers. Furthermore, we have performed idealized simulations to study the dynamical friction as a function of the self-interaction cross-section. In addition, we also studied how well cross-sections with different angular and velocity dependencies can be matched onto each other. Our main results can be summarized as follows:
@@
    The Knudsen number, Kn, is a good proxy for how much the self-interactions impact the dynamical friction, i.e., how close the system is to the collisionless or gaseous regime.
@@
    When the SIDM cross-section is large enough, $\mathrm{Kn}\lesssim 2$, we expect the dynamical friction in the supersonic regime to be significantly larger than in the collisionless case. Note, for typical binary BHs such a cross-section is many orders of magnitude smaller than current upper bounds on the self-interaction cross-section (Adhikari et al., 2022).
@@
The results we obtained are solely based on simulations of a constant density perturbed by a BH in a periodic box. This allowed us to gain insight into the qualitative behaviour but does not allow for accurately quantifying how strong the effect of self-interactions is on the dynamical friction force. In consequence, the next step would be to simulate a SIDM spike with an orbiting BH, similar to the work done by Kavanagh et al. (2020, 2024). Such a setup would allow us to precisely determine the strength of the dynamical friction force acting onto BHs orbiting in DM spikes when self-interactions are present. This is the subject of a forthcoming paper (Sabarish et al., in prep.).

### v2

@@ 4.2 Dependence on the strength of the self-interactions
Figure 5: Similar to Fig. 3, position and velocity of the BH particle as a function of time. The results are shown for several interaction strengths of the DM, indicated by the opacity of the lines. All simulations are for the large box with the high resolution of $N_{\mathrm{DM}}=5\times 10^{6}$ particles. For the simulation of gaseous matter, we used the same number of SPH particles.
@@
Given that the dynamical friction force differs between the collisionless and the gaseous case as described in Sect. 2, we expect that self-interactions can have an impact on the strength of the dynamical friction force. To study this in detail, we run simulations with SIDM. In Fig. 5, we show the results for fSIDM with different values for the momentum transfer cross-section. We also include the collisionless and gaseous case. It is visible that the self-interactions can enhance or reduce the dynamical friction, depending on the strength of the cross-section. For the gaseous regime, we expect the dynamical friction to be enhanced above the sound speed and reduced below. Accordingly, we find a larger velocity of the BH at the end of the simulation, when the BH velocity has fallen well below the sound speed of the DM.
@@
Figure 6: Surface density (top row) and the velocity dispersion (bottom row) for our test simulations at a time of $t=0.49\,\mathrm{yr}$. The panels display the different types of matter we investigate, with increasing collisionality from the left to the right. The density and one dimensional velocity dispersion are indicated following a logarithmic colour scale spanning a factor of two. For all panels of the same quantity, the same colour scale is used. In addition, we indicate the position of the BH particle with a red dot. We provide the time evolution of the surface density as a video available online.
@@
A detailed investigation of the dynamic friction force follows. But before, we take a look at the DM density perturbed by the BH. For the same simulations as shown in Fig. 5, we show the surface density of the DM in the top row of Fig. 6. Here, we can see that the shape and density of the wake depend on the strength of the self-interactions. We find that for $\sigma_{\mathrm{V}}/m\in\{3\times 10^{-8},3\times 10^{-7}\}\,(\mathrm{cm}^{2}\,\mathrm{g}^{-1})$ the densities in the wake are higher and at the same time the deceleration is stronger. For these small cross-sections, the scattering leads to more low-velocity particles in the density wake compared to CDM effectively enhancing the deceleration of the BH.
@@
In contrast, for the cases of even stronger self-interactions, we find that the densities in the wake are decreasing and the deceleration of the BH is weaker. In principle, this can be seen from Fig. 6, but we study this quantitatively in Sect. 4.3. We also note that the shape of the density wake changes for an increasing self-interaction strength, impacting the dynamical friction force as well. The reduced dynamical friction for the strong cross-sections is in line with the expectation of a weaker deceleration in the gaseous case for the subsonic regime.
@@
Figure 7: Dynamical friction force as a function of the Mach number for our simulations. This plot is analogue to Fig. 2, which gives the analytic description of the dynamical friction force. Different cross-sections are shown, indicated by the Knudsen number. Here, $\mathrm{Kn}=\infty$ corresponds to a collisionless medium and $\mathrm{Kn}=0.0$ to a gaseous medium. We note that in the simulation the BH initially starts at a high velocity, i.e. at the right side of the plot and slows down throughout the simulation, i.e. moves towards the left side of the plot.
For a more detailed analysis of the dynamical friction force and its dependence on the cross-section, we computed the force and show it as a function of the velocity (Fig. 7). The force is computed from the momentum of the BH and its change between consecutive snapshots. We note that the strongest SIDM cross-section shown is $\sigma_{\mathrm{V}}/m=3\times 10^{-6}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. Given our resolution, we are not able to faithfully simulate a cross-section as large as $\sigma_{\mathrm{V}}/m=3\times 10^{-5}\mathrm{cm}^{2}\mathrm{g}^{-1}$ because the mean free path becomes much smaller than the size of the kernel used for the self-interactions (see Appendix A and Fischer et al., 2024a).
We specify the strength of the cross-section in terms of the Knudsen number (Eq. 10). The collisionless case is indicated by $\mathrm{Kn}=\infty$ and the gaseous case by $\mathrm{Kn}=0.0$. In Fig. 7 we can see that the dynamical friction is initially increasing (starting at the highest velocity) and decreases for later times of the simulation. At the beginning of the simulation, where the BH moves with a supersonic velocity, the dynamical friction is the smallest in the collisionless case. At later stages, it depends on the cross-section whether the self-interactions enhance or reduce the deceleration of the BH. In Fig.7, we plot the dynamical friction as a function of the velocity. This shows that the decelerating force reaches its maximum roughly at the sound speed, as expected. Interestingly, the self-interacting run with an intermediate cross-section ($\mathrm{Kn}=1.35\times 10^{-1}$) shows the strongest dynamical friction, stronger than in the collisionless or gaseous regime. In principle, the figure is built analogously to Fig. 2, but a direct comparison is limited by the fact that our simulation setup may not fulfil the assumptions of the Ostriker formula for the dynamical friction in the gaseous case (see Sect. 2.2). This is for example illustrated by Fig. 3, where we can see that the exact results depend on the choice of the setup, in particular on the box size.
@@ 4.3 Properties of the wake
Figure 8: Density of the wake, i.e. the mean density about the point with the highest density, as a function of the velocity of the perturbing BH. The simulation results of the different cross-sections are indicated by their Knudsen number.
Figure 9: One-dimensional velocity dispersion of the wake, i.e. the velocity dispersion in the vicinity of the point with the highest density, as a function of the velocity of the perturber. We note that the velocity dispersion is computed after the bulk motion has been subtracted. The simulation results for different cross-sections are shown, indicated by the Knudsen number.
Figure 10: Mean velocity of the wake in the direction of the motion of the BH, i.e. the mean velocity in the vicinity of the point with the highest density. The results for different cross-sections are shown as a function of the velocity of the perturber. They are indicated by their corresponding Knudsen number.
Figure 11: Dark matter mass gravitational bound to the perturbing BH as a function of the velocity of the perturber. We indicate the different cross-sections by their Knudsen number.
To provide more insight into the increased dynamical friction for mildly collisional media, we computed further quantities such as the density of the wake, its velocity dispersion, and bulk motion as well as the DM mass gravitational bound to the BH.
@@
Furthermore, we show the velocity dispersion in the density wake in Fig. 9. We use the same volume as for the density computation. The velocity dispersion was computed after subtracting the mean velocity of the particles. The velocity dispersion in the wake increases compared to the ICs and is larger for stronger DM self-interactions. Interestingly, the cross-section with $\mathrm{Kn}=1.35\times 10^{-1}$, which gives the strongest dynamical friction, shows a substantially lower velocity dispersion compared to the larger cross-sections.
The bulk motion of the density wake is displayed in Fig. 10. The stronger the self-interactions the more efficiently the BH drags the DM along its trajectory, except for the largest cross-section for which the bulk velocity is again smaller. This is in line with the strength of the dynamical friction force found in Fig. 7, implying that when the DM is dragged along by the BH, it can more efficiently decelerate the BH.
Last, we computed the DM mass that is gravitationally bound to the BH. To do so, we compared the relative velocity of a DM particle and the BH to the escape velocity,
@@ 4.4 Angular dependence
The angular dependence of the self-interaction cross-section can in principle lead to qualitative differences in the evolution of astrophysical systems. This is for example well known for merging DM haloes (e.g. Fischer et al., 2023; Sabarish et al., 2024) or the abundance of satellites (Fischer et al., 2022). However, for systems that are rather relaxed, the angular dependence plays a minor role only and the effect of self-interactions with different angular dependencies can be mapped onto each other using the viscosity cross-section $\sigma_{\mathrm{V}}$ (Eq. 12, see also Yang & Yu, 2022).
To test whether this mapping also works for our setup to study dynamical friction, we ran simulations with a very anisotropic cross-section (using the fSIDM framework) and an isotropic cross-section. The cross-sections are chosen to have the same value in terms of $\sigma_{\mathrm{V}}/m$ or $\sigma_{\mathrm{T}}/m$ to test the matching with the momentum transfer cross-section (see Eq. 13) as well.
In Fig. 12, we compare the velocity of the BH for the different angular dependencies. We computed the ratio of the change in BH velocities for simulations with a matched isotropic cross-section to the fiducial fSIDM simulations. The two panels in Fig. 12 display the ratio of the velocities as a function of time for a weaker fSIDM cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ (top panel) and a stronger one of $\sigma_{\mathrm{V}}/m=3\times 10^{-7}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ (bottom panel). For a perfect matching, we would expect the ratio to be unity. As we can see, the viscosity cross-section (Eq. 12) provides a better match than the transfer cross-section (Eq. 13). This is the case for the weaker and stronger cross-section that we have simulated to test this.
Figure 12: Ratio between the velocity change, $\Delta v=v(0)-v(t)$, due to dynamical friction of an fSIDM cross-section and the transfer (blue) / viscosity (orange) matched isotropic cross-sections. The upper panel is for an fSIDM viscosity cross-section of $\sigma_{\mathrm{V}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ and the lower panel is for $\sigma_{\mathrm{V}}/m=3\times 10^{-7}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$.
@@ 4.5 Velocity dependence
The effective cross-section $\sigma_{\mathrm{eff}}$ (introduced by Yang & Yu, 2022) allows one to map cross-sections with different velocity dependencies onto each other. It integrates the cross-section with its velocity dependence over the assumed Maxwell-Boltzmann distribution of the scattering velocities, applying a weighting factor of $v^{5}$. The effective cross-section is given by
@@
We note that this is very similar to the matching procedure proposed by Yang et al. (2023); see also the work by Outmezguine et al. (2023). It has been shown that this works fairly well for isolated haloes and, in turn, velocity-dependent cross-sections can be approximately described with constant cross-sections. However, it is known that this does not work for systems that involve multiple velocity scales, such as merging galaxy clusters (Sabarish et al., 2024) or haloes containing substructure (Fischer et al., 2024b).
@@
Figure 13: Ratio of the velocity change, $\Delta v=v(0)-v(t)$, due to dynamical friction acting on the BH for a velocity-dependent and a velocity-independent (constant) cross-section. All cross-sections were matched using the effective cross-section (Eq. 11), i.e. all cross-sections correspond to $\sigma_{\mathrm{eff}}/m=3\times 10^{-8}\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. The upper panel displays runs with frequent self-interactions, and the lower panel is for isotropic scattering.
@@ 5 Discussion
For binary BHs, we find that the motion of the secondary BH through the DM spike would typically happen at Mach numbers larger than unity, in other words, in the supersonic regime. Moreover, the Mach number does only very slightly depend on the radius, as the mass of the DM spike in the relevant radial range is typically much smaller than the mass of the primary BH. Also, the spike index does not play a significant role in the motion of the secondary BH being supersonic. But more importantly, a non-zero eccentricity could lead to a subsonic phase about the apocentre of the orbit of the secondary BH. This implies that the supersonic regime where the dynamical friction is enhanced by self-interactions plays an important role and would lead to a faster decay of the orbit compared to a collisionless medium with the same density.
To study the dynamical friction with self-interactions being present, simulations are required. We used an idealised setup with a periodic box hosting initially a constant density. We have to note that this setup comes with drawbacks that limit the conclusion we can draw. In particular, it is difficult to create a stable setup where the box size is small enough such that the velocity dispersion of the DM counteracts the gravitational collapse. This effectively limits the maximal box size that can be studied, which also has an impact on the dynamical friction. This is because the box size is related to the maximum impact parameter $r_{\mathrm{max}}$ (see Eq. 7). Moreover, the setup does not always fulfil the assumptions of analytic formulas discussed in Sect. 2.2. Also, we cannot exclude that perturbations arising from the supersonic phase still affect the trajectory of the BH in the subsonic phase. Nevertheless, we are still able to make qualitative statements and gain new insights into the dynamical friction. Especially for this purpose, we tuned the setup to give a relatively larger impact on the BH velocity and make it easier to obtain qualitative differences being much larger than the numerical errors.
@@ 6 Conclusions
In this paper, we have reviewed analytic estimates for the dynamical friction force in the collisionless and gaseous regime, and discussed the two regimes in the context of binary BH mergers. Furthermore, we have performed idealised simulations to study the dynamical friction as a function of the self-interaction cross-section. In addition, we also studied how well cross-sections with different angular and velocity dependencies can be matched onto each other. Our main results can be summarised as follows:
@@
    The Knudsen number, Kn, is a good proxy for how much the self-interactions impact the dynamical friction, in other words, how close the system is to the collisionless or gaseous regime.
@@
    When the SIDM cross-section is large enough, $\mathrm{Kn}\lesssim 2$, we expect the dynamical friction in the supersonic regime to be significantly larger than in the collisionless case. We note that for typical binary BHs such a cross-section could be many orders of magnitude smaller than current upper bounds on the self-interaction cross-section (Adhikari et al., 2022).
@@
The results we obtained are solely based on simulations of a constant density perturbed by a BH in a periodic box. This allowed us to gain insight into the qualitative behaviour but does not allow for accurately quantifying how strong the effect of self-interactions is on the dynamical friction force. In consequence, the next step would be to simulate a SIDM spike with an orbiting BH, similar to the work done by Kavanagh et al. (2020, 2024). Such a setup would allow us to precisely determine the strength of the dynamical friction force acting onto BHs orbiting in DM spikes when self-interactions are present. This is the subject of a forthcoming paper (Sabarish et al., in prep).